# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

### v3

@@ IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [24, 22] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of numerical state tuples [27].
To bridge this gap, we transform each numerical state observation into a step-wise NL description and aggregate these into an NL trajectory observation. As shown in Fig. 1, NL description captures the semantic meaning of the state changes, such as “Move 5 meters West while facing North, the UAV moves to [0, 5]”. This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for code refinement.
Given the LLM-generated UAV operation code executed in simulation, Algorithm 2 presents the transformation from numerical state observations to the NL trajectory observation $O$. For each action identified in the $Code$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations. Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the difference between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $\Delta s=s_{\mathrm{current}}-s_{\mathrm{last}}$. For each element in the state difference $\Delta s$, the algorithm appends the corresponding NL description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$. The algorithm also appends the UAV’s NED position to $O$. Upon completion of all actions, the algorithm returns the complete NL trajectory observation $O$.
Input: $Code$: Code script
Output: $O$: Trajectory observation
@@
foreach  $action$ in $Code$  do
       $s_{\mathrm{last}}\leftarrow$ State Observation $(x,y,z,\theta)$
      Execute $action$
      if error executing $action$ then
             $O\leftarrow O$ + “Error in executing $action$ with error message $error$
@@
      $s_{\mathrm{current}}\leftarrow$ State Observation $(x,y,z,\theta)$
      $\Delta s\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
      foreach $element$ in $\Delta s$ do
             if $element$ in orientation then
                   $O\leftarrow O$ + Rotate $element$ degrees in Yaw.
@@
                  if $element$ in NED axis then
                         $O\leftarrow O$ + “Move $element$ meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
@@ IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the NL trajectory observation. The evaluator aims to identify deviations (if any) between the NL trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [40].
To improve the precision of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt. However, unlike the code generator, the evaluator does not produce code but instead generates evaluation feedback. To address this issue, we construct the evaluator’s system prompt with roles, rules, and references, as depicted in Fig. 3. Specifically, the roles define the LLM agent as an evaluator that compares the NL trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories. Additionally, the references are exemplar trajectories that serve as benchmarks to support the evaluation of NL trajectory observations. Furthermore, the system prompt requires the evaluation outcome to be informative and constructive by providing explanations about the depicted deviations. As demonstrated in Fig. 2, given a NL trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”). Consequently, the code generator is aware of the errors and could correctly refine the code in the subsequent refinement process.
Figure 3: The evaluation process begins by configuring the LLM agent (evaluator) via a system prompt framework that includes roles, rules, and references. Next, the trajectory observations and the task description are provided to the evaluator for comparison.
@@ IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 2. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new NL trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs ”YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
V Experiment
The goal of this experiment is to validate the effectiveness of our proposed LLM-driven closed-loop framework in enhancing the reliability of LLM-driven UAV operation code generation. To achieve this, the experiment is designed with three primary goals: (1) to evaluate the performance of our closed-loop feedback and refinement framework in LLM-driven UAV tasks; (2) to analyze the impact of different design choices on the performance of our framework; (3) to assess the behavior of our framework under diverse task scenarios.
@@ V-A Experimental Setup
V-A1 Environment
To ensure safe code execution without risking real-world failures, we conduct experiments in the “AirSim” simulator [41], which realistically simulates UAV (e.g., a quadcopter) dynamics and environmental interactions. During the experiment, UAV state information was fully accessible from the simulator in real time and utilized for closed-loop feedback and evaluation purposes.
V-A2 Tasks
We propose Basic and Advanced task sets to rigorously evaluate the performance of LLM-driven control code generation across varying levels of operational complexity [^1] The complete list of tasks is available at https://github.com/ai-uavsec/CLGSCE. The Basic set comprises 44 tasks adopted from GSCE [15]. Each task involves one or two intuitive discrete movements to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). In contrast, the Advanced set consists of 20 tasks that emulate real-world UAV deployment (e.g., “Take off and fly up 5 meters. You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left. Make sure the drone is oriented to the flying direction.”). These tasks demand reasoning about the objectives in the task description and require 6-19 UAV actions to reach the goal state. The advanced task set is intended to evaluate LLM’s multi-stage planning and complex spatial reasoning. Additionally, all tasks in both task sets were manually validated to avoid potential simulator-induced errors that could affect the evaluation process.
V-A3 Compared Methods
@@
Three recently proposed methods [15, 25, 24] are selected as the baselines for comparison with our approach, including
@@
  • GSCE: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement. [15]
@@
  • LLM Self-feedback (Self-Refine): A single LLM performs both code generation and self-refinement. The LLM analyzes NL trajectory observations to detect deviations from the task and refines code based on its self-assessment.  [25]
@@
  • Feedback with Numerical State Observations (Numerical): This method introduces a secondary LLM serving as evaluator designed according to section IV-D. The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement. [24]
All methods described above employ “o3-mini” (o3-mini-2025-01-31) [29] as the foundational LLM. The generated code was then deployed on a quadcopter in AirSim [41] to evaluate the performance of each method. To mitigate the randomness of LLM’s output [42], we repeat each task three times in the experiment.
V-B Evaluation Setup
V-B1 Metrics
The performance is evaluated using two metrics as GSCE [15]: Completeness and Success Rate (SR).
@@ V-B1 Metrics
Completeness quantifies the proportion of goal states successfully reached, calculated as the ratio of the total count of correct actions to the total number of required actions of the ground truth trajectory in each task. The completeness offering insights into performance across the intermediate control process.
$Completeness=\frac{\sum\text{correct actions}}{\mathrm{len}(l)}$ (1)
@@
SR is grounded in real-world operational requirements, wherein the UAV must fly the trajectory precisely as humans instructed. As formalized in Eq. 2, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory.
$Success=\begin{cases}1,&Completeness=1,\\ 0,&Completeness<1.\end{cases}$ (2)
V-B2 Ground Truth
@@
The ground truth is represented by a list of state transitions. It is a vector of four elements: $[x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation. We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
@@ V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table I. For the Basic tasks, all methods showed remarkable performance. Specifically, the GSCE, Self-Refine, and Ours methods achieved $100\%$ in both SR and completeness, while the Numerical achieved $95.5\%$ SR and $97.3\%$ completeness. This demonstrates their ability to generate correct code for a few-stage planning and basic spatial reasoning. However, the slightly lower performance of the Numerical method suggests that relying on numerical state observations can occasionally lead to misinterpretation in evaluation, resulting in incomplete or incorrect refinements.
Performance on the Advanced tasks declines notably across all methods as a result of increased task complexity. The open-loop GSCE method achieves only $66.7\%$ SR, indicating that refinement becomes critical as task complexity increases. The Self-Refine method performs worse, with $50.0\%$ SR and $74.3\%$ completeness, proving that relying on a single LLM for both generation and self-evaluation amplifies bias [40], resulting in introducing additional errors in refinement. The Numerical method achieves $73.3\%$ SR and $92.4\%$ completeness, showing that it can detect some deviations; however, its numeric perspective still fails to provide reliable evaluation feedback for refinement. In contrast, our proposed method substantially outperforms all baselines, achieving $85.0\%$ SR and $98.5\%$ completeness on Advanced tasks. The overall results demonstrate that our proposed framework significantly enhances the reliability of LLM-driven UAV operation code generation, particularly for complex tasks that require multi-stage planning and complex spatial reasoning.
TABLE I: Average SR and Completeness of Proposed Methods
@@
Base Task Advanced Task SR Completeness SR Completeness GSCE [15] $100.0\%$ $100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [25] $100.0\%$ $100.0\%$ $50.0\%$ $74.3\%$ Numerical [24] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $100.0\%$ $100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
V-C2 Failure Cases Analysis
We further analyze the failure cases (cases that $Success=0$) in the Advanced task set. The analysis focuses on three aspects: (1) the number of failures and average completeness for each failed task ID with the respective ground truth action length (Fig. 4); (2) the average completeness of each failed task ID (Fig. 5); (3) the amount of task that experienced failure and the average completeness of all failed tasks (Table II).
The results in Fig. 4 reveal a clear trend that all methods exhibit a notable increase in failures as task complexity increases (i.e., more likely to fail on tasks with longer action length). Specifically, the baseline GSCE method failed in $11$ out of $20$ tasks ($55\%$ of the Advanced set), with an average completeness of $77.4\%$. This indicates that the GSCE method can reason through a substantial portion of the task correctly, however, it still struggles to maintain full alignment with the objectives of the task. This result highlights the need for improvement in the completeness of failed tasks to reduce overall failure rates. The Self-Refine method exhibited the weakest performance, with failures on $17$ tasks ($85\%$ of Advanced tasks). Moreover, the failures in Self-Refine were the least complete among all methods, demonstrating that performing evaluation and refinement on a single LLM not only fails to correct errors but may also introduce more errors. The Numerical method failed on $9$ tasks ($45\%$ of Advanced tasks) and achieved an average completeness of $69.5\%$ only. These results indicate that numerical state observations are insufficient for reliable code refinement and may even lead to erroneous refinement.
Figure 4: Failure analysis across the four methods on the Advanced task set. The x-axis represents the task IDs that are sorted in ascending order based on the length of the task’s action steps. And the y-axis presents two metrics for each task: (1) the count of failures observed over three repetitions (left y-axis); (2) the length of the ground truth action steps (right y-axis).
Figure 5: Average completeness for each failed task on the Advanced task set over four methods. The smaller the value, the less complete the failed task is.
In contrast, our method achieves the best performance, failures occurred only on $4$ tasks ($20\%$ of Advanced tasks), and the completeness of failed tasks arose at $91.6\%$. Moreover, the completeness in Fig. 5 demonstrates that even when failures occur, our method produces code that remains closely aligned with task objectives. These findings underscore the effectiveness of our method in refining errors and improving the completeness of LLM-driven UAV operation, which contributes to an increased success rate in handling complex UAV operation tasks.
TABLE II: Amount of Task ID Experienced Failure and Average Completeness for All Failed Tasks
Task ID Amount $\downarrow$ Failed Tasks Completeness $\uparrow$
GSCE [15] $11$ $77.4\%$
Self-Refine [25] $17$ $48.5\%$
Numerical [24] $9$ $69.5\%$
Ours $\mathbf{4}$ $\mathbf{91.6\%}$
V-C3 Number of Refinement Iterations
[…]

### v4

@@ IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [7, 8] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [16].
To bridge this gap, we transform each numerical state observation into a step-wise semantic description and aggregate these into a semantic trajectory observation. As an example shown in Fig. 4, the semantic description captures the semantic meaning of the state changes and UAV position, such as “Move 5 meters East while facing North, the UAV moves to [0, 5, -11.5]”. This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for guiding the code generator on code refinement.
Algorithm 2 presents the transformation from numerical state observations to the semantic trajectory observation $O$. For each action identified in the $l$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations $x,y,z,\theta]$ and then transformed to semantic trajectory descriptions as the example illustrated in Fig. 4. Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the state transition $T$ between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $T=s_{\mathrm{current}}-s_{\mathrm{last}}$. For each element in the state transition $T$, the algorithm appends the corresponding semantic description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$. The algorithm also appends the UAV’s NED position vector to $O$ for embedding the positional information. Upon completion of all actions of $l$, the algorithm returns the complete semantic trajectory observation $O$.
Figure 4: An example of transforming numerical state observations into semantic trajectory observation. UAV state vector of $x,y,z,\theta]$ denotes (North, East, Down, Yaw).
Input: $l$: Code script
Output: $O$: Semantic trajectory observation
@@
foreach  actionaction in $l$ do
 $s_{\mathrm{last}}\leftarrow$ Numerical state observation $x,y,z,\theta]$
  Execute actionaction
 if error executing actionaction then
    $O\leftarrow O$ + “Error in executing actionaction with error message errorerror
@@
 $s_{\mathrm{current}}\leftarrow$ Numerical state observation $x^{\prime},y^{\prime},z^{\prime},\theta^{\prime}]$
 $T\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
 foreach elementelement in $T$ do
    if elementelement in orientation then
       $O\leftarrow O$ + Rotate elementelement degrees in Yaw.
@@
       if elementelement in NED axis then
          $O\leftarrow O$ + “Move elementelement meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
@@ IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the semantic trajectory observation. The evaluator aims to identify deviations (if any) between the semantic trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [18].
To improve the accuracy of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt. However, unlike the code generator, the evaluator does not produce code but instead generates evaluation feedback. To address this issue, we construct the evaluator’s system prompt with roles, rules, and references, as depicted in Fig. 5. Specifically, the roles define the LLM agent as an evaluator that compares the semantic trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories. Additionally, the references are example trajectories that serve as benchmarks to support the evaluation of semantic trajectory observations. Furthermore, the system prompt requires the evaluation outcome to be informative and constructive by providing explanations about the depicted deviations. As demonstrated in Fig. 3, given a semantic trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”). Consequently, the code generator is aware of the errors and could correctly refine the code in the subsequent refinement process.
Figure 5: The evaluation process begins by configuring the LLM agent (evaluator) via a system prompt framework that includes roles, rules, and references. Next, the semantic observations and the task description are provided to the evaluator for comparison.
@@ IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 3. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new semantic trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs “YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
V Experiment and Evaluation
@@ V-A Experimental Setup
In our experiment, the implementation of our proposed framework and comparison methods [5, 10, 7] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [20] as the foundational LLM. Based on the “o3-mini”, two separate LLMs are configured as the code generator and evaluator in our framework. The widely used UAV simulator, “AirSim” [27], is tailored to support the simulation component of our framework for code execution and state observations. The UAV model selected for experiment and evaluation is the default quadcopter in AirSim. The detailed implementation of our framework and experiment is available at our project website  [^1] Project Website: https://github.com/ai-uavsec/CLGSCE.
UAV Task Dataset: In our experiment, two categories of UAV tasks are selected to evaluate the performance of our framework under different operational complexities, i.e., Basic tasks and Advanced tasks. Each Basic task involves one or two intuitive actions to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). We use the Basic task dataset established by existing research [5], which contains 44 tasks in total. Differently, the Advanced tasks in our experiment require complex reasoning about the UAV’s position and orientation dynamics and typically require 6-19 actions to reach the goal state. We established the Advanced task dataset by emulating real-world UAV operation tasks, which consists of 20 tasks with different complexities. As summarized in Table I, there are four levels of complexity: “Low”, “Medium”, “High”, and “Very High”. The “Low” tasks involve flying simple geometric patterns. The “Medium” tasks introduce orientation requirements while flying the patterns (e.g., always orienting to the flying direction). The “High” tasks require flying complex patterns with orientation requirements (e.g., fly a figure 8 while orienting in the flying direction). The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flies the path reversely). The complete list of both Basic and Advanced tasks is available at our project website \@footnotemark All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
TABLE I: Examples of Tasks in Advanced Task Dataset
Complexity Num. of Tasks Avg. Steps Example Tasks
Low 3 9 “You are going to scan square areas. First, take off and climb up 5 m. Fly a 5 m square (forward/right/back/left). Then shift 5 m south, and fly the same square pattern again.”
Medium 9 11 “You are going to view the borders of square areas. Please take off and fly up 5 meters. You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left in world axis. Make sure the drone is oriented to the flying direction.”
High 4 17.5 “You are going to fly a figure of 8. Please take off and fly up 5 meters. Fly a figure of 8 on a flat, horizontal plane with each side of 5 meters. The left square is on your left-rear side, and the right square is on your right-front side. You should begin with the left square by flying left. The right square should start from moving forward. Make sure the drone is oriented in the flying direction.”
Very High 4 17.5 “You are going to examine a cube. Take off and fly up 5 meters. You will examine a square area. You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left. To examine the square area, the drone should orientate perpendicular to the moving direction on each side of the square. Next, ascend another 5 meters and fly the square pattern in reverse order to examine the same area.”
@@
Baseline Methods for Comparison: Three recently proposed methods [5, 10, 7] are selected as the baselines for comparison with our proposed approach, including
@@
  • GSCE [5]: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement.
@@
  • LLM Self-feedback (Self-Refine) [10]: A single LLM performs both code generation and self-refinement. The LLM analyzes semantic trajectory observations to detect deviations from the task and refines code based on its self-assessment.
@@
  • Feedback with Numerical State Observations (Numerical) [7]: This method introduces a secondary LLM serving as evaluator designed according to section IV-D. The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement.
V-B Evaluation Metrics
In our evaluation, two key metrics Completeness and Success Rate (SR) are used for performance measurement. Given a UAV task, the completeness measures the ratio of actions required by the task that are completed successfully, and the SR measures whether the entire task is completed successfully. The detailed definition of these two metrics, as well as the ground truth used in our evaluation, is provided below.
@@ V-B Evaluation Metrics
Completeness: As formalized in Eq. 1, completeness for task $i$ computed as the ratio between the count of actions correctly executed ($a^{\mathrm{correct}}_{i}$) and the total number of actions in the task’s ground truth action sequence ($l^{\mathrm{gt}}_{i}$). This metric offers insights into performance across the intermediate control process, where a higher completeness indicates a greater proportion of the task that has been successfully completed.
$Task\_Completeness_{i}=\frac{|a^{\mathrm{correct}}_{i}|}{|l^{\mathrm{gt}}_{i}|}$ (1)
Given $n$ tasks, the completeness is calculated as:
$Completeness=\frac{1}{n}\sum_{i=1}^{n}Task\_Completeness_{i}$ (2)
@@
SR: As formalized in Eq. 3, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory. Then we compute the SR across $n$ tasks as Eq. 4. This metric is grounded in real-world operational requirements, wherein the UAV must perform the task precisely as instructed.
$Success_{i}=\begin{cases}1,&Task\_Completeness_{i}=1,\\ 0,&Task\_Completeness_{i}<1.\end{cases}$ (3)
$SR=\frac{1}{n}\sum_{i=1}^{n}Success_{i}$ (4)
@@
Ground Truth: The ground truth is represented by a list of state transitions. It is a vector of four elements: $x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation. We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
@@ V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table II. The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [28].
TABLE II: SR and Completeness of Proposed Methods
@@
Basic Task Advanced Task SR Completeness SR Completeness GSCE [5] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [10] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [7] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
For the Basic tasks, all methods showed remarkable performance. It is notable that the Numerical method [7] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness. This indicates that relying on numerical state observations can lead to misinterpretation in evaluation, even when handling simple tasks.
When moving to the Advanced tasks that have a higher complexity, our method still achieved a reliable performance with an $85\%$ SR and a $98.5\%$ completeness. This result shows that $98.5\%$ of required actions in all evaluated tasks are completed correctly, and $85\%$ of tasks are entirely completed without any error. As a comparison, the reliability of baseline methods drops notably when the task complexity increases due to the limitations caused by factors such as amplified bias, limited numerical reasoning capabilities of LLMs, and open-loop design without refinement. While the impact of these factors is relatively negligible when handling basic tasks, their limitations are amplified when the task complexity increases, and affect the logical reasoning capabilities of LLMs to generate operation code correctly.
Figure 6: Failure counts of different methods for each advanced task at different complexity levels
Figure 7: Completeness of each failed task of different methods at different complexity levels
TABLE III: Failure Cases Completeness in Each Level of Complexity
Low Medium High Very Overall
High
GSCE [5] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
Self-Refine [10] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
Numerical [7] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
Ours $\mathbf{90.0\%}$ $\mathbf{94.4\%}$ $\mathbf{90.4\%}$ $\mathbf{91.6\%}$
Failure Case Analysis: To uncover the detailed performance of each method in its failed advanced tasks, we performed an in-depth failure case analysis. In particular, we aim to understand how many actions are correctly completed when a task fails. As summarized in Table III, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur. Differently, the average completeness of the compared baseline methods ranges from $48.5\%$ to $77.4\%$ for their failed tasks, which indicates errors occur for a significant portion of the required actions in these tasks. Fig. 6 further details the failure count on each failed task of all methods in our three rounds of experiment with task complexity and length of actions information. Our method has failures on only four tasks during three rounds of experiments. As a comparison, the baseline methods failed 11 (GSCE [5]), 17 (Self-Refine [10]), and 9 (Numerical [7]) tasks, respectively. In addition, Table III and Fig. 7 also show that our method has a stable high completeness of $90\%+$ for all failed tasks at different complexity levels. In contrast, the baseline methods have not only significantly lower average completeness ($21\%$ to $90.9\%$) on failed tasks at different levels but also a large variation in performance of different tasks.
Based on our evaluation results, our method achieves a higher reliability compared with the baseline methods, not only because of its high confidence in generating fully correct operation code, but also its high completeness of the required actions of tasks when a small amount of errors occur.
TABLE IV: Evaluator Configurations Description and Example System Prompts
Evaluator Configuration Description
(1) Roles This configuration defines the role for the LLM as the code evaluator, and its goal is to evaluate whether the UAV flight trajectory is aligned with the task description and output standardized feedback for the code generator.
(2) Roles + Rules Besides the roles in (1), this configuration also includes specific rules for the LLM to follow. The rules introduce explicit evaluation criteria on the UAV’s actions, spatial positions, flight paths, coordinate system, and offset tolerance that direct the LLM’s evaluation of the UAV trajectory.
(3) Roles + References Besides the roles in (1), this configuration also provides reference examples to support evaluation and generate feedback. These references include pairs of task descriptions and corresponding semantic trajectory observations.
(4) Roles + Rules + References This configuration combines roles, rules, and references to provide comprehensive support information for the LLM to perform its job correctly.
Example System Prompts
Roles 1. You will reason about the difference between the task description and drone actions.
2. If the task description and drone actions are aligned, you should output “YES” only. Otherwise, you should output “NO”; please output where the drone actions mismatch the task description if your answer is “NO”.
Rules 1. For tasks that require perpendicularly examining a square area, you must check whether the drone is oriented perpendicularly inwards the square interior; facing toward the exterior is not allowed.
2. You should reason about the spatial locations of square patterns from actions and ensure the drone flies the square pattern in the correct spatial position as outlined in the task description.
3. The drone uses the NED coordinate system (world coordinates) represented as [x, y, z], the positive X axis is North/forward, the positive Y axis is East/right, and the negative Z axis is up/climb.
References Here is a correct path of a 4-meter square pattern (forward/right/backward/left) for reference: 1. Fly 4 meters North in the world axis while facing north. The drone moves to [4, 0]. 2. Fly 4 meters East in the world axis while facing north. The drone moves to [4, 4]. 3. Fly 4 meters South in the world axis while facing north. The drone moves to [0, 4]. 4. Fly 4 meters West in the world axis while facing north. The drone moves to [0, 0].
*Note: more examples of system prompts for roles, rules, and references are available at the project website: https://github.com/ai-uavsec/CLGSCE.
V-C2 Number of Refinement Iterations
In our proposed framework, the operation code generated by the Code Generator LLM will be refined through multiple iterations. The increased rounds of iterations provide additional opportunities to improve the code quality, however, they also introduce additional system cost (e.g., LLM API cost and processing time). To balance this tradeoff and identify an appropriate number of iterations, we performed experiments of our framework on advanced tasks with 0 to $10$ rounds of iterations. Our results presented in Fig. 8 show the changes of SR and completeness with the different number of iterations. Specifically, both SR and completeness are boosted after the iteration is introduced for refinement, and then have a slowly increasing trend until the $6\textsuperscript{th}$ iteration. After that, the performance of the subsequent iterations becomes relatively stable in terms of both SR and completeness, and even declines slightly. This is because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [29, 30], and thus may perform unnecessary changes after sufficient refinement has been performed. Our detailed examination identified that such a situation occurred on task 7 in our experiment. This observation demonstrates that iterative refinement can effectively contribute to the LLM-driven UAV operation code generation. Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff. According to our evaluation results, we set the maximum number of iterations to $6$. However, this number may need to be adjusted for different types of tasks. For example, setting a lower number when the tasks to complete have a relatively low complexity.